Determine the following probabilities:

X         -2          -1         0           1           2

f(x)    0.125     0.25   0 .25     0.25      0.125

PX<=2)

P(X> -2)

P(-1<= X,<= 1)

   P(X<- -1)
Answer #1

0.125 P(x2) = P(x=-2) Þ(xy-2)- 1 - P(x5-2) = 1-0.0125 = 10.875 P(-14X51) = P(x=-1) +P(x=0) + P(x-1) = 0.25+0.25 +0.25 - P(x-1